Essay 2: List the different modes of dichloro methane. Which modes are IR active and which are Raman active? (20 points)

as detailed as possible with drawings please. thank you!
Answer #1

Answer: Dichiono methane (CH2012) cí o point group of chelin. a car ен. С12 Total number of vibrational mode = 3 n-6 = 15-6 =7. C-ci asymmetric Stretching Cur@ -> IR Achte my Raman active 8. H-c-H in plane Rocking (bending) @r) o~ -IR Active > Raman